x^2+16x=0

I have to complete the square. I thought I needed to take way the x^2 and put on the other side and then I am left with 16x instead of 16. I could square 16 which is 4. But I don't know what to do with the other x?

Answer by stanbon(75887)

complete the square.
x^2+16x=0
---
x^2 + 16x + 8^2 = 8^2
----
(x+8)^2 = 64
----
(x+8)^2-64 = 0
Factof:
[(x+8)-8][(x+8)-8] = 0
----
x+8-8 = 0 or x+8+8 = 0
x = 0 or x = -16
======================
Cheers,
Stan H.